All of them, please

What is the minimum number of bits you need to represent the number (500)10 in binary ?


1- What is the problem when we represent the number (325)10 as (11 01 101)2 ?

2- What is the minimum number of bits you need to represent the number (500)10 in binary ?

3- Writing (325)10 in BCD we will get (011010101)2


4- Write (2014)10 in BCD ?
Answer #1

Answer :

To represent decimal 500 in binary , number of bits required are : log (500) / log(2) , where log is log base 10. = 9 bits.

1) as given binary number = 2^6 + 2^5 + 2^ 3 + 2^2 + 2^ 0 = 64+32+8+4 +1 = 109 so it is not equal to 325.

2) as shown above , 9 bits.

3) in bcd : 3 = 011

2 = 010

5 = 101

So 325b= 011010101, so given statement is true.

6 ) 2 = 010

1 = 001

0 = 000

4 = 0100

So 2014 = 0100000010100